# AgentArch: A Comprehensive Benchmark to Evaluate Agent Architectures in Enterprise

### v1

@@ A.1 Use Cases and Set Up
Within each use case, we represent a wide array of potential edge cases. For example, in the Requesting Time Off use case we have user requests which should be rejected and some that should be accepted. Within the rejected category, we include requests that should be rejected for a wide variety of reasons such as invalid leave types, conflicting leaves, and most significantly, lack of leave balance. For lack of balance, we make sure to include complex cases where models must correctly count the number of days across multiple months or take into account leap years and the differing number of days in different months.
For each sample in a use case, the ground truth values for expected tool inputs, expected outcome, and expected tool order were human annotated and verified.
@@
A.2 Thinking Tools Examples
  1. 1.
    Synthesizing information
    tool: synthesize_collected_information
    args:
    synthesized_information (str):
    result (same as input):
  2. 2.
    Math
    tool: math
    args:
    input (str):
    result (same as input)::
A.3 Prompts
Prompts were designed to be simple and representative examples of both traditional function calling-style prompts and ReAct prompts. For function calling prompts, the tools and directions are placed in the system prompt while the user prompt includes the main objective and conversation history thus far. We also include a note to always call a tool (since communication with the user is not supported), always use real input values (since some models are prone to using placeholders instead), and to select the finish tool when finished since this is critical for control flow as well as getting a self reported summary of each agents actions which is passed to other agents in the summarized memory management style. We attempt to convey the same information in the ReAct style prompt however as is typical with ReAct prompting, all information is in the system prompt and a JSON response is required. The ReAct prompts include a sample JSON output format that asks the model for ”thought” (reasoning) and an ”action” and ”observation.”.
A.4 Additional Metrics

### v2

@@ Appendix A Use Cases and Set Up
Within each use case, we represent a wide array of potential edge cases. For example, in the Requesting Time Off use case we have user requests that should be rejected and some that should be accepted. Within the rejected category, we include requests that should be rejected for a wide variety of reasons such as invalid leave types, conflicting leaves, and most significantly, lack of leave balance. For lack of balance, we make sure to include complex cases where models must correctly count the number of days across multiple months or take into account leap years and the differing number of days in different months.
For each sample in a use case, the ground truth values for expected tool inputs, expected outcome, and expected tool order were human annotated and verified. We construct the use cases in such a way that a limited set of deterministic ground truths options can be determined for every data sample.
Appendix B Thinking Tools Examples
  1. 1.
    Synthesizing information
    tool: synthesize_collected_information
    args:
    synthesized_information (str):
    result (same as input):
  2. 2.
    Math
    tool: math
    args:
    input (str):
    result (same as input)::
Appendix C AI Usage
We use AI in this paper for generating the code required for running the benchmark, for proof reading and paraphrasing, and for creating the graphics and figure formatting.
Appendix D Prompts
Prompts were designed to be simple and representative examples of both traditional function calling-style prompts and ReAct prompts. For function calling prompts, the tools and directions are placed in the system prompt while the user prompt includes the main objective and conversation history thus far. We also include a note to always call a tool (since communication with the user is not supported), always use real input values (since some models are prone to using placeholders instead), and to select the finish tool when finished since this is critical for control flow as well as getting a self reported summary of each agents actions which is passed to other agents in the summarized memory management style. We attempt to convey the same information in the ReAct style prompt however as is typical with ReAct prompting, all information is in the system prompt and a JSON response is required. The ReAct prompts include a sample JSON output format that asks the model for "thought" (reasoning) and an "action" and "observation.".
Appendix E Additional Metrics